OpenStudy (anonymous):

@kevinzhengli Checking account A charges a monthly service fee of $20 and a wire transfer fee of $3, while checking account B charges a monthly service fee of $30 and a wire transfer fee of $2. How many transfers would a person have to have for the two accounts to cost the same? A. 31 B. 10 C. 0 D. 21 @mattt9

OpenStudy (anonymous):

what do you think the first step is?

OpenStudy (anonymous):

20+3x = 30+2x x = 10 B.

OpenStudy (anonymous):

nice.

OpenStudy (anonymous):

i thought we had to subtract something?